# Simple homotopy types of even dimensional manifolds

### v2

@@ Abstract.
Given a closed $n$-manifold, we consider the set of simple homotopy types of $n$-manifolds within its homotopy type, called its simple homotopy manifold set. We characterise it in terms of algebraic $K$-theory, the surgery obstruction map, and the homotopy automorphisms of the manifold. We use this to construct the first examples, for all $n\geq 4$ even, of closed $n$-manifolds that are homotopy equivalent but not simple homotopy equivalent. In fact, we construct infinite families with pairwise the same properties, and our examples can be taken to be smooth for $n\geq 6$.
Our key examples are homotopy equivalent to the product of a circle and a lens space. We analyse the simple homotopy manifold sets of these manifolds, determining exactly when they are trivial, finite, and infinite, and investigating their asymptotic behaviour. The proofs involve integral representation theory and class numbers of cyclotomic fields. We also compare with the relation of $h$-cobordism, and produce similar detailed quantitative descriptions of the manifold sets that arise.
@@ 1.1. Background and main results
The homeomorphism classification made use of Reidemeister torsion to distinguish homotopy equivalent lens spaces. While trying to understand this more deeply, J.H.C. Whitehead defined the notion of simple homotopy equivalence [Whitehead-simplicial-spaces, Whitehead-incidence, Whitehead-combinatorial, Whitehead-she]. A homotopy equivalence $f\colon X\to Y$ between CW complexes is said to be simple if it is homotopic to the composition of a sequence of elementary expansions and collapses; see Section 2 for the precise definition. Chapman [Ch74] showed that every homeomorphism $f\colon X\to Y$ between compact CW complexes is a simple homotopy equivalence, and so we have:
[…]
@@
This article constructs the first examples of closed manifolds, in all even dimensions $2k\geq 4$, that are homotopy but not simple homotopy equivalent. In fact, we produce infinite families.
From now on, $n\geq 4$ will be an integer and an $n$-manifold will be a compact, connected, $\operatorname{CAT}$ $n$-manifold where $\operatorname{CAT}\in\{\operatorname{Diff},\operatorname{PL},\operatorname{TOP}\}$ is the category of either smooth, piecewise linear, or topological manifolds. Manifolds will also be assumed closed unless otherwise specified. We will also frequently make use of the following hypothesis.
@@
If $n=4$, we assume $\operatorname{CAT}=\operatorname{TOP}$ and restrict to manifolds $M$ such that $\pi_{1}(M)$ is good in the sense of Freedman (see [FQ, Freedman-book-goodgroups]).
Theorem A.
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Then there exists an infinite collection of orientable $\operatorname{CAT}$ $n$-manifolds that are all homotopy equivalent but are pairwise not simple homotopy equivalent.
It is currently open whether all topological 4-manifolds are homeomorphic to CW complexes. Thus, the definition of simple homotopy equivalence given previously does not apply when $n=4$ and $\operatorname{CAT}=\operatorname{TOP}$. For a more general definition that does apply in this case, see Definition 2.3.
We note that if $M_{1}\simeq N_{1}$ and $M_{2}\simeq N_{2}$ are two pairs of homotopy equivalent (but not necessarily simple homotopy equivalent) odd dimensional manifolds, then $M_{1}\times M_{2}\simeq_{s}N_{1}\times N_{2}$ (Corollary 2.33), so even dimensional examples cannot be constructed in such a straightforward way from odd dimensional examples.
1.2. Simple homotopy manifold sets
Beyond finding the first examples of even-dimensional manifolds that are homotopy equivalent but not simple homotopy equivalent, in order to quantify this phenomenon we study the following manifold sets. Writing $\simeq$ for homotopy equivalence and  $\simeq_{s}$ for simple homotopy equivalence, for a $\operatorname{CAT}$ $n$-manifold $M$ we introduce the simple homotopy manifold set
@@ 1.2. Simple homotopy manifold sets
$\mathcal{M}^{h}_{s}(M)\coloneqq\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\simeq M\}/\simeq_{s}$
This is the set of manifolds homotopy equivalent to $M$ up to simple homotopy equivalence. Our main examples will be manifolds homotopy equivalent to $S^{1}\times L$, where $L$ is a lens space.
We prove that for manifolds of the form $M=S^{1}\times L$, the size of $\mathcal{M}^{h}_{s}(M)$ can be trivial, finite and arbitrarily large, and infinite, and we determine precisely when each eventuality occurs. In particular the existence of the latter case implies Theorem A. For an integer $m\geq 2$, let $C_{m}$ denote the cyclic group of order $m$, and let $C_{\infty}$ denote the infinite cyclic group.
@@
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $M^{n}_{m}=S^{1}\times L$, where $L$ is an $(n-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, $m\geq 2$. Then
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|$ only depends on $n$ and $m$, and is independent of the choice of $L$ or $\operatorname{CAT}$;
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$;
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free;

### v3

@@ Abstract.
Given a closed $n$-manifold, we consider the set of simple homotopy types of $n$-manifolds within its homotopy type, called its simple homotopy manifold set. We characterise it in terms of algebraic K-theory, the surgery obstruction map, and the homotopy automorphisms of the manifold. We use this to construct the first examples, for all $n\geq 4$ even, of closed $n$-manifolds that are homotopy equivalent but not simple homotopy equivalent. In fact, we construct infinite families of manifolds that are all homotopy equivalent but pairwise not simple homotopy equivalent, and our examples can be taken to be smooth for $n\geq 6$.
Our examples are homotopy equivalent to the product of a circle and a lens space. We analyse the simple homotopy manifold sets of these manifolds, determining exactly when they are trivial, finite, or infinite, and investigating their asymptotic behaviour. The proofs involve integral representation theory and class numbers of cyclotomic fields. We also compare with the relation of $h$-cobordism, and produce similar detailed quantitative descriptions of the manifold sets that arise.
@@ 1.1. Background and main results
The homeomorphism classification made use of Reidemeister torsion to distinguish homotopy equivalent lens spaces. While trying to understand this more deeply, J.H.C. Whitehead defined the notion of simple homotopy equivalence [Whitehead-simplicial-spaces, Whitehead-incidence, Whitehead-combinatorial, Whitehead-she]. A homotopy equivalence $f\colon X\to Y$ between CW complexes is said to be simple if it is homotopic to the composition of a sequence of elementary expansions and collapses. He also defined the Whitehead group [Whitehead-she] $\operatorname{Wh}(G)$ of a group $G$, and the Whitehead torsion $\tau(f)\in\operatorname{Wh}(\pi_{1}(Y))$. He proved the fundamental result that $\tau(f)=0$ if and only if $f$ is simple. See Section 2 for the precise definition.
Chapman [Ch74] showed that every homeomorphism $f\colon X\to Y$ between compact CW complexes is a simple homotopy equivalence, and so we have:
$\mbox{homeomorphism}\Rightarrow\mbox{simple homotopy equivalence}\Rightarrow\mbox{homotopy equivalence}.$
@@
This article constructs the first examples of closed manifolds, in all even dimensions $2k\geq 4$, that are homotopy but not simple homotopy equivalent. In fact, we produce infinite families. Writing $\simeq$ for homotopy equivalence and $\simeq_{\mathrm{s}}$ for simple homotopy equivalence, we note that if $M_{1}\simeq N_{1}$ and $M_{2}\simeq N_{2}$ are two pairs of homotopy equivalent (but not necessarily simple homotopy equivalent) odd dimensional manifolds, then $M_{1}\times M_{2}\simeq_{\mathrm{s}}N_{1}\times N_{2}$ (Corollary 2.33), so even dimensional examples cannot be constructed in such a straightforward way from odd dimensional examples.
Theorem A.
Let $n\geq 4$ be even. Then there exists an infinite collection of closed, connected, orientable $n$-manifolds that are all homotopy equivalent but are pairwise not simple homotopy equivalent. If $n>4$, these manifolds can be taken to be smooth, and if $n=4$, they are topological.
Our examples are manifolds homotopy equivalent to $S^{1}\times L$, for certain lens spaces $L$.
It is currently open whether all topological 4-manifolds are homeomorphic to CW complexes. Thus, the definition of simple homotopy equivalence given previously does not apply. For a more general definition that applies in this case, see Definition 2.3.
1.2. Simple homotopy manifold sets
From now on, $n\geq 4$ will be an integer and all manifolds will be assumed closed and connected. In order to quantify the difference between homotopy and simple homotopy equivalence, for a $\operatorname{CAT}$ $n$-manifold $M$ we introduce the simple homotopy manifold set
$\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M):=\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\simeq M\}/\simeq_{\mathrm{s}}$
where $\operatorname{CAT}\in\{\operatorname{Diff},\operatorname{PL},\operatorname{TOP}\}$ is the category of either smooth, piecewise linear, or topological manifolds. This is the set of manifolds homotopy equivalent to $M$ up to simple homotopy equivalence. The term “manifold set” is used in contrast to structure sets: an element of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ is represented by manifolds without a chosen homotopy equivalence to $M$.
We will frequently restrict to the case of $\operatorname{CAT}$ $n$-manifolds where $\operatorname{CAT}$ satisfies the following hypothesis which depends on $n$.
@@ 1.2. Simple homotopy manifold sets
If $n>4$, we take $\operatorname{CAT}\in\{\operatorname{Diff},\operatorname{PL},\operatorname{TOP}\}$. If $n=4$, we take $\operatorname{CAT}$ to be the full subcategory of $\operatorname{TOP}$ with objects $4$-manifolds $M$ such that $\pi_{1}(M)$ is good in the sense of Freedman.
Remark 1.2.
For good groups, the surgery sequence for 4-manifolds is defined and exact, and the $s$-cobordism theorem is available (see Section 1.2). The class of good groups contains finite groups and solvable groups, and is closed under subgroups, quotients, extensions and direct limits. For more details and references see [FQ, Freedman-book-goodgroups]. We have $\pi_{1}(S^{1}\times L)\cong C_{\infty}\times C_{m}$ where $C_{\infty}$ is infinite cyclic and $C_{m}$ is cyclic of order $m\geq 2$. Since $C_{\infty}\times C_{m}$ is good, 1.1 applies in the case of Theorem A.
To study simple homotopy manifold sets, we will make use of a further equivalence relation, $h$-cobordism. A cobordism $(W;M,N)$ of closed manifolds is an $h$-cobordism if the inclusion maps $M\to W$ and $N\to W$ are homotopy equivalences. The notion of an $h$-cobordism is central to manifold topology. Smale’s $h$-cobordism theorem [smale-pe, Smale-h-cob, Milnor-h-cob], together with its extensions to other categories and dimension 4 in [Stallings-Tata, Kirby-Siebenmann:1977-1, FQ], states that under 1.1, every simply-connected $h$-cobordism is $\operatorname{CAT}$-equivalent to the product $M\times I$. To generalise the $h$-cobordism theorem to non-simply connected manifolds, one also considers $s$-cobordisms. An $s$-cobordism $(W;M,N)$ is a cobordism for which the inclusions $M\to W$ and $N\to W$ are simple homotopy equivalences. The $\operatorname{CAT}$ $s$-cobordism theorem [Barden, Mazur:1963-1, Stallings-Tata, Kirby-Siebenmann:1977-1, FQ] states that under 1.1, $\operatorname{CAT}$ equivalence classes of $h$-cobordisms based on $M$ are in bijection with $\operatorname{Wh}(\pi_{1}(M))$, with the bijection given by taking the Whitehead torsion of the inclusion $M\to W$ (Theorem 3.1). In particular every $s$-cobordism is a product. This theorem underpins manifold classification in dimension at least $4$.
The $s$-cobordism theorem is one of the tools that we can use to construct homotopy equivalent but not simple homotopy equivalent manifolds (see Sections 3.1 and 4.2), and the examples constructed this way are always $h$-cobordant. This leads us to consider two refined versions of the problem of finding homotopy but not simple homotopy equivalent manifolds, where the manifolds are also required to be $h$-cobordant, or required not to be $h$-cobordant. We introduce the corresponding variations of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$:
@@
$\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ $=\left\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\text{ is $h$-cobordant to }M\right\}/\simeq_{\mathrm{s}}$
$\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$ $=\left\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\simeq M\right\}/\langle\simeq_{\mathrm{s}},\operatorname{hCob}\rangle$
where $\langle\simeq_{\mathrm{s}},\operatorname{hCob}\rangle$ denotes the equivalence relation generated by simple homotopy equivalence and $h$-cobordism. These sets arise naturally in the computation of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ in Theorem B below.
In Theorem B below, we characterise the simple homotopy manifold sets $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$, $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$, and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$, for a $\operatorname{CAT}$ $n$-manifold $M$, in terms of the Whitehead group, the homotopy automorphisms of $M$, and the surgery obstruction map. We then apply these characterisations to the manifolds $S^{1}\times L$ and compute the objects that appear in Theorem B for this case.
We first introduce notation which appears in the theorem statement. An orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$ determines an involution $x\mapsto\overline{x}$ on the Whitehead group $\operatorname{Wh}(G)$ (see Section 2.2), and we write $\operatorname{Wh}(G,w)$ to specify that $\operatorname{Wh}(G)$ is equipped with this involution. If $w\equiv 1$ is the trivial character, then we omit it from the notation. Define:
$\mathcal{J}_{n}(G,w)$ $=\left\{y\in\operatorname{Wh}(G,w)\mid y=-(-1)^{n}\overline{y}\right\}\leq\operatorname{Wh}(G,w),$
$\mathcal{I}_{n}(G,w)$ $=\left\{x-(-1)^{n}\overline{x}\mid x\in\operatorname{Wh}(G,w)\right\}\leq\mathcal{J}_{n}(G,w).$
The Tate cohomology group $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is canonically identified with $\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ (see Proposition 9.4), and we denote the quotient map by
Next let
$\varrho\colon\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\rightarrow L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)$
be the map from the Ranicki-Rothenberg exact sequence (3.2) (see [Shaneson-GxZ], [Ranicki-80-I, §9]).
For $M$ a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w$, let
$\sigma_{\mathrm{s}}\colon\mathcal{N}(M)\rightarrow L_{n}^{\mathrm{s}}(\mathbb{Z}G,w)$
be the surgery obstruction map (see Section 3.2). The homotopy automorphisms $\operatorname{hAut}(M)$ of $M$ is the group of homotopy classes of self-homotopy equivalences of $M$. There is an action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$ (as a set) such that if $f\colon N\rightarrow M$ is a homotopy equivalence and $g\in\operatorname{hAut}(M)$, then $\tau(f)^{g}=\tau(g\circ f)$. Denote the quotient map by
$q\colon\operatorname{Wh}(G,w)\rightarrow\operatorname{Wh}(G,w)/\operatorname{hAut}(M).$
The subset $\mathcal{J}_{n}(G,w)$ is invariant under the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$, and there is an induced action on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. The subsets
$(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\subseteq\mathcal{J}_{n}(G,w)\quad\text{ and }\quad\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
are both invariant under the induced actions of $\operatorname{hAut}(M)$. In general $\mathcal{I}_{n}(G,w)$ is not invariant under the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$, but when it is, then $q(\mathcal{I}_{n}(G,w))=\mathcal{I}_{n}(G,w)/\operatorname{hAut}(M)$.
The following theorem is the basis of our main results (see Theorems 4.11 and 4.16).
@@
Let $M$ be a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, with $\operatorname{CAT}$ as in 1.1. There is a commutative diagram
${\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${q(\mathcal{I}_{n}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})/\operatorname{hAut}(M)}$
It follows from the exact sequence (3.2) that the image of the homomorphism $\psi\colon L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is $\varrho^{-1}(\{0\})$, hence it is contained in $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$, and so we obtain the following corollary.
Corollary 1.3.
Let $M$ be a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, with $\operatorname{CAT}$ as in 1.1. If $\psi\colon L_{n+1}^{\mathrm{h}}(\mathbb{Z}G,w)\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is surjective, then we have
$\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)\cong\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M),\quad\quad\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M).$
The advantage of Corollary 1.3 is that $\mathcal{J}_{n}(G,w)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ (and, more generally, $\operatorname{Im}\psi$) only depend on $G$ and $w$, while $\operatorname{Im}\sigma_{\mathrm{s}}$ depends on $M$ a priori. This allows us to apply Corollary 1.3 by separately analysing the involution on $\operatorname{Wh}(G,w)$ and the homotopy automorphisms of $M$. In cases where $M$ can be regarded as a manifold in multiple categories, the sets $q(\mathcal{I}_{n}(G,w))$, $\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)$ do not depend on the choice of $\operatorname{CAT}$.
We can also use Theorem B to describe when one of the simple homotopy manifold sets is nontrivial, see Propositions 4.20, 4.21, and 4.22. We introduce the following notation:
$T(M)$ $=\left\{\tau(g)\mid g\in\operatorname{hAut}(M)\right\}\subseteq\mathcal{J}_{n}(G,w)$
$U(M)$ $=\left\{\pi(\tau(g))\mid g\in\operatorname{hAut}(M)\right\}\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
Corollary 1.4.
Let $M$ be a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, with $\operatorname{CAT}$ as in 1.1. Then
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)|>1$ if and only if $\mathcal{I}_{n}(G,w)\setminus T(M)$ is nonempty.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)|>1$ if and only if $(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus T(M)$ is nonempty.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)|>1$ if and only if $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})\setminus U(M)$ is nonempty.
To prove Theorem B, first we observe that if $N$ is a manifold homotopy equivalent to $M$, then the set of Whitehead torsions of all possible homotopy equivalences $f\colon N\to M$ forms an orbit of the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$. By considering the different restrictions and equivalence relations on these manifolds $N$, we obtain maps from the various simple homotopy manifold sets of $M$ to subsets/subquotients of $\operatorname{Wh}(G,w)/\operatorname{hAut}(M)$. We then verify that these maps are injective, and the remaining task is to determine their images.
For $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$, we use the $s$-cobordism theorem. For every $x\in\operatorname{Wh}(G,w)$ there is an $h$-cobordism $(W;M,N)$ with Whitehead torsion $x$, and then $\tau(f)=-x+(-1)^{n}\overline{x}$ for the induced homotopy equivalence $f\colon N\rightarrow M$ (see Proposition 2.38). This shows that every element of $\mathcal{I}_{n}(G,w)$ can be realised as the torsion of a homotopy equivalence $N\rightarrow M$ for some $N$ that is $h$-cobordant to $M$. For $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$, we need to describe the set of values of $\pi(\tau(f))$ for all homotopy equivalences $f\colon N\to M$. This set is the image of a natural map $\widehat{\tau}\colon\mathcal{S}^{\mathrm{h}}(M)\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ from the homotopy structure set of $M$. We show that $\widehat{\tau}$ fits into a commutative diagram involving the surgery exact sequence and the Ranicki-Rothenberg exact sequence (3.2) (see Proposition 3.7), and use that diagram to show that its image is $\varrho^{-1}(\operatorname{Im}\sigma_{\mathrm{s}})$. Finally, the characterisation of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ is obtained by combining the results on $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$.
1.3. Computations of simple homotopy manifold sets
We prove that for manifolds of the form $M=S^{1}\times L$, the size of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ can be trivial, finite and arbitrarily large, or infinite, and we determine precisely when each eventuality occurs. In particular the existence of the latter case implies Theorem A. For an integer $m\geq 2$, let $C_{m}$ denote the cyclic group of order $m$, and let $C_{\infty}$ denote the infinite cyclic group.
Theorem C.
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $L$ be an (n1)(n-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, for some $m\geq 2$, and let $M^{n}_{m}=S^{1}\times L$. Then
  1. (a)
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|$ only depends on $n$ and $m$, and is independent of the choice of $L$ or $\operatorname{CAT}$;
  2. (b)
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$;
  3. (c)
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free;